random samples of size 4 are drawn with replacement from a finite population 3,6,9


  1. How many possible samples of size 4 are possible?
  2. Find the mean of the sample means.
  3. Find the variance of the sample.
  4. Find the standard deviation of the sample.
1
Expert's answer
2021-11-17T17:53:06-0500

1.

We have population values "3,6,9" population size "N=3" and sample size "n=4." Thus, the number of possible samples which can be drawn without replacement is


"N^n=3^4=81"

2.

In sampling with replacement the mean of all sample means equals the mean of the population:


"\mu_{\bar{X}}=\mu=\dfrac{3+6+9}{3}=6"


3.

When sampling with replacement the variance of all sample means equals the variance of the population divided by the sample size


"\sigma^2=\dfrac{1}{3}((3-6)^2+(6-6)^2+(9-6)^2=6"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}=\dfrac{6}{4}=1.5"



4.


"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{\dfrac{\sigma^2}{n}}=\dfrac{\sigma}{\sqrt{n}}"




"=\dfrac{\sqrt{6}}{\sqrt{4}}=\sqrt{1.5}\approx1.224745"